RECENT CHANGES IN THE SPFIT PROGRAM SET

H. M. PICKETT, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109.

In this talk the current capabilities of the SPFIT/SPCAT program set will be described. New additional capabilities include: (1) a new option that allows minimization of the rms width of a blend, (2) provision for alternative phase factors for Wang basis functions, (3) improved capabilities for ℓ-doubled states, (4) capabilities for internal rotation effects on energies and intensities.

The discussion period after the talk has been extended to 45 minutes to answer questions and to deal with specific SPFIT applications.